Reports 1-1 of 1 Clear search Modify search
AdV-ISC (Commissioning up to first full interferometer lock)
pinto, vardaro, ruggi - 20:05 Monday 24 January 2022 (54592) Print this report
ISC shift: robustness of lock at CARM zero, new signal for OS monitoring and test for the BS AA loop

Task of the shift was to further study the robustness of the lock at CARM zero. Two main tasks were performed:

1 - test the procedure developed in the past shift in order to minimize the optical spring and to reach a good alignment working point;
2 - implement an improved signal to monitor the behaviour of the optical spring wrt to the alignment conditions.

Despite a couple of locks during which we unlocked during the engage of the SSFS (8.29.02 UTC, 15.02.29 UTC) and one during the handoff of CARM to MC (9.49 UTC), we obtain the first relevant lock at CARM null 3f at 10.07.10 UTC. The action performed during this lock are the following:

10.08.14 UTC: engage of the SSFS and handoff of the longitudinal loops to the 1f signals;

This time we didn't close immediately the dither loops, but we first performed the alignment procedure in order to minimize the fringe, with the final aim to engage them with zero offsets, once we obtained an optical spring close to zero.
Thus, we started to inject noise on DARM at 10.10 UTC to monitor the DARM TF with respect to the usual alignment procedure to reach the target values (DARM phase at 100 HZ = 0.8 rad, pole of TF at 14 Hz). Such procedure produced the expected effect of reducing the optical spring. In details, between 10.12 UTC and 11.10 UTC we moved the SR in the TX (+) and TY (-) directions of an overall misalignment of respectively 2 urad and 0.6 rad. Moreover, we adjusted the longitudinal offset of SRCL up to a final value of 6.5. During these movements, we were iteratively adjusting the BS in TX (-) in order to minimize the fringe.

N.B. Since at 10.34 we seemed to have reached a good alignment working point (in terms of optical spring), we closed the dither loops with 0 offsets, but after a while (11.39) we added small offsets and proceeded with the aligment.

At 11.36 we performed a first test of closing the BS AA loop, exploiting the B8_DC signal demodulated with the BS dithering signals. The loop stayed closed for few minutes but at 11.39.48 UTC the ITF unlocked due to big flashes on B1p_DC.

Other than monitoring the DARM TF, we were looking, as a figure of merit, at the OS monitoring signal calibrated in Hz(LSC_OS_LF_mag_Hz) which showed as a global trend, the expected behaviour of going towards 0, as we were minimizing the spring. However such signal, since it is a magnitude of the DARM signal, has no sign and also, below 6 Hz doesn't give any meaningfull information (see 54585).

For these reasons, for the rest of the shift we worked on a new signal, which is created from the demodulation of B7_DC with the DARM line at 74.4 Hz. The channel name is LSC_NArm_B7_OS_I and has been implemented on Acl. This new signal has not been tested yet, but it will during the afternoon shift.

We left the interferometer ready for the enxt ISC activities.


Images attached to this report
Search Help